ou have been managing a $5 million portfolio that has a beta of 1.05 and a required rate of return of 12.925%. The current risk-free rate is 4%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 0.75, what will be the required return on your $5.5 million portfolio? Do not round intermediate calculations. Round your answer to two decimal places.'

?
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

12.925=4+1.05*(market rate-4)

(12.925-4)=1.05*(market rate-4)

market rate=(12.925-4)/1.05+4

=12.5%

Hence Required rate for $500,000=4+0.75*(12.5-4)

=10.375%

Hence required rate for portfolio=Respective return*Respective weight

=(5,000,000/5,500,000*12.925)+(500,000/5,500,000*10.375)

=12.69%(Approx).